OpenStudy (anonymous):

1/4+7/12+2/3 sum using or difference by using benchmarks

OpenStudy (anonymous):

I do not know what benchmarks are, but I can solve this if you want me too...but it might not be the correct way that you need. Do you want me to show you how I do it ?

OpenStudy (anonymous):

you have to make them have the same denominators when adding or subtracting fractions. The common denominator of 4,12,and 3 is 12. 1/4 = 3/12 7/12 = 7/12 2/3 = 8/12 so now we have : 3/12 + 7/12 + 8/12 = 18/12 which reduces to 3/2 or in mixed number form = 1 1/2